OpenStudy (anonymous):

chemist dean dalton is given a clump of metal and is told that it is sodium. He finds that the metal weighs 145.5g. he places it into a nonreactice liquid in a square prism whose base measures 10cm on each edge. If the metal is indeed sodium, how high should the liquid level rise? The density of sodium is .97g/cm3 (answer and work please)

OpenStudy (toxicsugar22):

hi math mind

OpenStudy (anonymous):

haha hello sugar (:

OpenStudy (toxicsugar22):

can u help me after u help her

OpenStudy (anonymous):

sure. I'll be right there

OpenStudy (anonymous):

can I have help....?

OpenStudy (anonymous):

sorry, i though you had left

OpenStudy (anonymous):

yeah me too haha

OpenStudy (anonymous):

answer please

OpenStudy (anonymous):

yup, working on it

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[Density = {(mass) \over (volume)}\]\[Vsquare = (width)*(length)*(height)\]where: s = side/edge The initial volume of the square prism is: 10^3 = 10*10*10 = 1000 cm^3 to find the final volume (volume after the metal is added) we can use the density formula:\[(0.97) = {145.5 \over x}\](0.97)x = 145.5\[x = {145.5 \over 0.97} = 150 cm^3\] Therefore, if (volume) = (width)*(length)*(height) and if you add something the height would increase by x so, (10+x)\[150 cm^3 = 10*10*(10+x)\]150 = 100(10+x) 150 = 1000 + 100x 100x = 150-1000 x = 850/100 x = 8.5 cm rise